# Open-Vocabulary SAM3D: Understand Any 3D Scene

### v2

@@ Abstract
Open-vocabulary 3D scene understanding presents a significant challenge in the field. Recent advancements have sought to transfer knowledge embedded in vision-language models from the 2D domain to 3D domain. However, these approaches often require learning prior knowledge from specific 3D scene datasets, which limits their applicability in open-world scenarios. The Segment Anything Model (SAM) has demonstrated remarkable zero-shot segmentation capabilities, prompting us to investigate its potential for comprehending 3D scenes without the need for training. In this paper, we introduce OV-SAM3D, a universal framework for open-vocabulary 3D scene understanding. This framework is designed to perform understanding tasks for any 3D scene without requiring prior knowledge of the scene. Specifically, our method is composed of two key sub-modules: First, we initiate the process by generating superpoints as the initial 3D prompts and refine these prompts using segment masks derived from SAM. Moreover, we then integrate a specially designed overlapping score table with open tags from the Recognize Anything Model (RAM) to produce final 3D instances with open-world label. Empirical evaluations conducted on the ScanNet200 and nuScenes datasets demonstrate that our approach surpasses existing open-vocabulary methods in unknown open-world environments. Full codes and models will be made publicly available.
@@ 1 Introduction
Recently, research on foundation models has been booming [10, 4, 28, 27, 29, 13, 1, 39], and they have shown emergent capabilities of complex reasoning, reasoning with knowledge, and out-of-distribution robustness [38]. Beyond the field of NLP, foundation models profoundly influence many other fields [25, 21, 37], especially the emergence of Segment Anything Model (SAM) [21] which has brought a new breakthrough in the computer vision community. The zero-shot segmentation performance of SAM is amazing, so some works attempt to extend the powerful generalization ability to other downstream tasks [30, 24, 7, 43], and 3D scene segmentation and understanding is a significant transferable domain [41, 40, 36].
Subject to the difficulty and cost of collecting 3D data, transferring knowledge of foundation models from the 2D domain to the 3D domain is a feasible solution to achieve zero-shot learning in the 3D scene. Existing studies such as SAM3D [41] and SAMPro3D [40] can effectively segment 3D instances, but these methods are limited by the fact that SAM only achieves segmentation without recognition, unable to achieve 3D scene understanding.
Considering the challenge of open-vocabulary 3D scene understanding, the model should locate and recognize objects in the given 3D scene through the text guidance regardless of whether the objects have been seen before. Related works such as OpenScene [26] and OpenMask3D [36] can achieve open-vocabulary 3D scene understanding given the premise of having studied similar scenes in a dataset, but they exhibit poor performance in domain transfer. For the purpose of understanding any 3D scene, our research adequately explore the powerful zero-shot segmentation ability of SAM without the need for 3D encoders pre-training on specific datasets.
@@
In this work, we introduce OV-SAM3D, a universal open-vocabulary 3D scene understanding framework which can complete understanding tasks on any given scene without prior knowledge about the scene. Inspired by the traditional graph-based segmentation algorithm [14], our OV-SAM3D introduces superpoints as initial 3D prompts, then revises these 3D prompts to rough 3D masks and calculates the overlapping score through the back-projected segmentation results of SAM [21]. Through the recognition of RAM [44] employed in posed images of the given 3D scene and the screening of ChatGPT, OV-SAM3D can get open-world image tags and matches these tags with initial 3D prompts by vision-language model such as CLIP [27]. Based on the overlapping score and matched tags, OV-SAM3D refines the rough 3D masks to obtain the final 3D instances with labels. Compared with the existing works, our method not only breaks through the limitation of class-agnostic masks when transferring SAM knowledge from 2D to 3D domain, but also achieves better open understanding results than previous open-vocabulary 3D scene understanding algorithms(Fig. 1). As a universal 3D scene understanding algorithm, OV-SAM3D can understand arbitrary 3D scene ignorant of prior knowledge about the scene.
@@
In order to compare with existing methods, we conduct experiments on the well-studied ScanNet200 dataset [31], but the closed-set environment goes against our original intention of understanding any 3D scene in an unknown open-world. Therefore, we further conduct experiments on the outdoor nuScenes dataset [5] where existing algorithms are rarely applied, and the results show that OV-SAM3D is superior to existing state-of-the-art methods in the open-world.
Figure 1: Paradigm comparison with contemporary methods. (a): SAM3D [41] successfully transfer SAM knowledge to 3D field, but it can only generate the class-agnostic segmentation results. (b): OpenMask3D [36] implements the open-vocabulary scene understanding task, but it requires training a 3D proposal network under supervision and the queried text comes from the specific dataset or human knowledge. (c): Our OV-SAM3D can not only transfer the rich knowledge of SAM to 3D field, but also further understand any 3D scene without training through the combination of multiple foundation models.
@@
  • We present OV-SAM3D, a universal framework for open-vocabulary 3D scene understanding, capable of interpreting arbitrary 3D scenes without the need for any prior knowledge.
@@ 2 Related Work
Foundation model is clearly defined to highlight its role in the field of artificial intelligence [3]. The foundation models first shone brightly in the field of NLP with the emergence of transformer structures [10, 4, 13, 34, 33]. Currently, not only in the field of NLP, multimodal foundation models concentrating all data information from various modalities are also in full swing [27, 15, 12, 15]. CLIP [27] conducts contrastive learning through a large number of image-text pairs, enabling the model to understand the semantic relationship between images and texts, which can be used for various visual and language tasks. Based on this, MaskCLIP [45] extends the image-text classification ability through feature upsampling to coarse image segmentation ability. Tag2Text [19] introduces tagging tasks which obtain image tags from image-text pairs to guide the visual-language model learning visual-language features better. RAM [44] innovatively combines tagging tasks and captioning tasks which expands the recognition range of tagging. As a landmark foundation model in the vision community, SAM [21] can generate high-quality masks through point prompts, box prompts, and text prompts. Our research attempts to effectively combine the recognition capability of RAM, the localization capability of SAM and the classification capability of CLIP to complete more complex scene understanding tasks.
Zero-Shot 3D Scene Segmentation.
@@
The impressive performance of SAM in zero-shot image segmentation tasks has greatly promoted the development of foundation models in computer vision. Its derivative models can be applied to fields with high difficulty in image segmentation such as medical images, remote sensing images and even extended to 3D point clouds to achieve zero-shot 3D scene segmentation [7, 8, 30, 24, 6, 17]. Recent researches relies on SAM to implement foundation model functions in the 3D domain because of the enormous cost of collecting massive 3D point cloud data that meets the training needs of foundation models. SAM3D [41] segments multiple posed images of the given 3D scene through the panoramic segmentation function of SAM, then projects the segmentation mask to the 3D scene and iteratively merges adjacent point clouds with the bidirectional merging approach until capturing the 3D mask of the entire scene. After the preliminary attempts with SAM3D, SAMPro3D [40] samples points from the given scene and projects them into the posed images as point prompts assisting SAM to achieve more accurate segmentation. SAI3D [42] introduces the over-segmented results as super points, and calculates affinity to merge the super points to obtain the final 3D segmentation masks. These zero-shot 3D scene segmentation methods inherit the class-agnostic characteristic from SAM, which cannot recognize the segmented instance. Compared with these methods, our research challenges open-vocabulary 3D scene understanding tasks by the union of multiple foundation models.
Open-Vocabulary 3D Scene Understanding.
@@
Given any ambient text query, open-vocabulary 3D scene understanding tasks [11, 18] aim to locate and identify matched instances, regardless of whether the model has seen these instances before. OpenScene [26] trains a 3D encoder learning to extract dense features of the point cloud in the 3D scene through a distillation loss with the frozen 2D Scene Understanding model [16, 23], and ensembles with features extracted from the posed images to CLIP embeddings for an arbitrary set of queries. OpenIns3D [20] firstly uses a trained 3D network to learn class-agnostic mask proposals, and then projects these proposals onto synthetic scene-level images [2, 22]. With the 2D scene understanding model detecting the position and category of objects in the scene-level images, OpenIns3D matches the projected mask proposals to achieve open-vocabulary 3D scene understanding. OpenMask3D [36] also utilizes Mask3D [35] to generate class-agnostic masks for 3D scenes, which are projected onto the posed image of the given scene as point prompts. Then it extracts average CLIP features of multi-scale image crops obtained from segmentation results by SAM, which are matched with the CLIP embeddings of the query texts to achieve 3D scene understanding. Compared to previous studies, the significant improvement of our method is that we abandon training-required 3D encoders which limit the generalization ability of models and explore the collaborative effects of foundation models to achieve a universal open-vocabulary 3D scene understanding algorithm.
@@ 3 Methodology
Fig. 2 shows the detailed framework of our OV-SAM3D. Given any scene and its corresponding set of posed RGB-D images, inspired by recent works [41, 32, 42], we generate approximately over-segmentation superpoints. The core of OV-SAM3D lies in the merging these superpoints into accurate 3D instances and obtaining corresponding labels during the process. Firstly, we select high-quality superpoints as the initial 3D prompts and project them onto posed images as point prompts guiding SAM to fine segment. The segmented mask is projected back into the 3D scene for revising the initial 3D prompts to coarse masks and extract the overlapping scores of the coarse masks. Meanwhile, we utilize RAM [44] to identify open tags from the corresponding frames and filter out unreasonable tags through ChatGPT. Combining the overlapping scores and matched open tags with CLIP, we can obtain the final 3D instances with label.
Figure 2: Overview of our OV-SAM3D that consists of two sub-modules: 1) SAM-centric Coarse Mask Generation (Sec. 3.1) first adapts graph-based over-segmentation method to generate superpoints and selects some as initial 3D prompts to guide SAM. Then we revise the initial 3D prompts through the masks of SAM and create an overlapping score table. 2) Open Tags Guided Coarse Mask Merging (Sec. 3.2) combines the overlapping score table and reasonable open tags recognized by RAM, thus we can achieve the open-vocabulary 3D scene understanding task to get 3D instances with label.
@@ 3.1 SAM-centric Coarse Mask Generation
Directly processing unstructured 3D point clouds in 3D scenes significantly influences computational efficiency, while grouping 3D points with similar geometric properties into continuous regions as superpoints remarkably accelerates the processing of OV-SAM3D. Referring to the graph based image segmentation algorithm [14], we obtain $M$ superpoints $\{S_{m}\in\mathbb{R}^{P_{m}\times 3}\}_{m=1}^{M}$ and select high-quality $N$ superpoints as the initial 3D prompts $\{Q_{n}\}_{n=1}^{N}\subseteq S$ based on $P_{i}$, which is the number of point cloud included in $S_{m}$. Then, we project the 3D prompts onto the top $T$ views with the lowest degree of occlusion, and sample the projected pixel points as point prompts to guide SAM for fine segmentation.
@@
Considering the initial 3D prompts as the origin of the final 3D instance, we design an overlapping score table $F\in\mathbb{R}^{M\times N}$ to determine which origin all superpoints belong to. Then we back-project the $n$-th 2D mask on the $t$-th view as $BP_{nt}\in\mathbb{R}^{P_{nt}\times 3},n\leq N,t\leq T$ for revising the randomly segmented 3D mask to a coarse 3D mask, where $P_{nt}$ represents the number of points including in the 3D back-projection.
The basis of this revision process and subsequent merging procedure is the designed overlapping score table $F$ which is generated following Algorithm 1. Firstly, we calculate the overlap region of other superpoints with the back-projection in the same 3D scene. Then, we add the score at the corresponding position $F_{ij}$ which depends on whether the overlap region of $\{S_{m}\cap Q_{n}\}$ meets a certain threshold $\theta$. After all superpoints are merged into the corresponding 3D prompts, we can get an overlapping score table $F$, as visualized in Fig. 4. For the overlapping score table, along the row axis, the maximum index in each line $F_{m}^{1\times N}$ represents which 3D prompts the superpoints $S_{m}$ should belong to, and then we merge all superpoints into the corresponding 3D prompts to obtain a coarse 3D masks $R_{c}\in\mathbb{R}^{P\times N}$ where $P$ is the total number of point cloud in the given scene. Following the column axis, each column $F_{n}^{M\times 1}$ can represent the distribution of overlap between each coarse 3D mask and all superpoints, which we define as the overlapping score of this coarse 3D mask.
Figure 3: Visualization of overlapping features. We calculate the similarity between the overlapping scores of the coarse mask pairwise. If the similarity is higher than a certain threshold, it indicates that these two coarse masks belong to the same instance. Here, we visualize some overlapping score similarity of the queried coarse masks with all masks and color points from high similarity to low similarity.
@@ 3.2 Open Tags Guided Coarse Mask Merging
Due to the uncertainty of initial 3D prompts in any given 3D scene, the coarse 3D masks need to be merged again for generating the final high-quality 3D instance masks. We remerge the coarse 3D masks primarily relying on the judgment whether the total score $TS_{n}\in\mathbb{R}^{M\times 1}$ of the overlapping scores between coarse 3D masks exceed a certain threshold which is derived from the overlapping score itself and user selected parameters:
@@
where $TS_{n}$ is the similarity score of $n$-th coarse 3D mask with other masks. $\tau$ is a threshold that can be modified by user-interaction. Note that OV-SAM3D delivers some user-adjustable hyper-parameters adapting to any given 3D scene in a open world. The threshold $\tau$ here is one of them, and we just limit its range, specific values can be adjusted in different 3D scenes for better understanding.
Limited by the narrow perspective of posed images, there may be some large instances with no overlap regions in different views, which cannot be directly merged through overlapping score. Therefore, we adopt an updatable merging strategy, which continuously updates the coarse 3D mask to the merged result and records the composition of merged coarse 3D masks during the merging process. Through the above algorithm, we can ultimately obtain class-agnostic 3D instances with composition records.
@@
OV-SAM3D implements a universal open 3D scene understanding framework, and our strict and practical requirement is that the model cannot learn any prior knowledge about pending understanding 3D scenes in advance. Without knowing what objects may exist in a 3D scene, we introduce RAM to perform preliminary recognition on the corresponding posed images of the scene as shown in Fig. 5, roughly identifying which instances or labels are present in the 3D scene. However, in the open tags library of RAM, there are not only instance tags, but also some tags that represent colors, material properties, room types and so on. These instance independent tags are not conducive to 3D scene understanding tasks. Therefore, we add ChatGPT following RAM as a filter to select the instance tags recognized by RAM and generate the final open instance tags.
@@
After merging $N$ coarse masks into $K$ 3D instances and recognizing open instance tags, we match the tags with the class-agnostic 3D instance segmentation masks through CLIP to achieve the absolute open-vocabulary 3D scene understanding task. In order to ensure the accuracy of matching without increasing algorithm complexity, we refuse to project the 3D mask onto posed images again. Instead, we extract image crops features from existing SAM segmented 2D masks through CLIP’s image encoder. Then match the image crops features $\{FC_{k}\in\mathbb{R}^{I_{k}\times 768}\}_{k=1}^{K}$ (where $I_{k}$ is the numble of course masks contained in k-th instance according to the composition records mentioned in 3.2) with the CLIP text features $FT\in\mathbb{R}^{C\times 768}$ extracted from open tags by cosine similarity $CS_{ki}$:
@@
where $CS_{ki}$ represents the similarity of course mask between $C$ text queries. For each coarse mask, we can choose the tag with the highest similarity as the label $\{Label_{i}\}_{i=1}^{I_{k}}$, but the matching result cannot be directly used as the label of the final 3D instance. We list the coarse 3D masks contained in each 3D instance based on the composition records, and align the open tags $Label_{i}$ with the highest feature similarity $CS_{ki}$ among them as the final label $Label_{k}$ for corresponding 3D segmentation instance. Subsequent ablation study in Tab. 3 shows slight differences in the selection of the highest feature similarity and the highest matching number and the above selection can get better results.

### v3

# Open-Vocabulary SAM3D: Towards Training-free Open-Vocabulary 3D Scene Understanding

@@ Abstract
Open-vocabulary 3D scene understanding presents a significant challenge in the field. Recent works have sought to transfer knowledge embedded in vision-language models from 2D to 3D domains. However, these approaches often require prior knowledge from specific 3D scene datasets, limiting their applicability in open-world scenarios. The Segment Anything Model (SAM) has demonstrated remarkable zero-shot segmentation capabilities, prompting us to investigate its potential for comprehending 3D scenes without training. In this paper, we introduce OV-SAM3D, a training-free method that contains a universal framework for understanding open-vocabulary 3D scenes. This framework is designed to perform understanding tasks for any 3D scene without requiring prior knowledge of the scene. Specifically, our method is composed of two key sub-modules: First, we initiate the process by generating superpoints as the initial 3D prompts and refine these prompts using segment masks derived from SAM. Moreover, we then integrate a specially designed overlapping score table with open tags from the Recognize Anything Model (RAM) to produce final 3D instances with open-world labels. Empirical evaluations on the ScanNet200 and nuScenes datasets demonstrate that our approach surpasses existing open-vocabulary methods in unknown open-world environments. Full codes and models will be made publicly available.
@@ 1 Introduction
Recently, research on foundation models has been booming [9, 3, 30, 29, 31, 12, 1, 40], and they have shown emergent capabilities of complex reasoning, reasoning with knowledge, and out-of-distribution robustness [39]. Foundation models have profoundly influenced various fields beyond NLP [26, 20, 38], particularly with the advent of the Segment Anything Model (SAM) [20], which has marked a significant breakthrough in the computer vision community. The zero-shot segmentation performance of SAM is remarkable, prompting some efforts to extend its powerful generalization capabilities to other downstream tasks [32, 25, 6, 44]. One major transferable domain for such advancements is 3D scene segmentation and its further understanding [42, 41, 43, 37].
Figure 1: Paradigm comparison with contemporary methods. (a): SAM3D [42] leverages SAM to achieve zero-shot 3D segmentation but only generates the class-agnostic segmentation results. (b): OpenMask3D [37] implements open-vocabulary scene understanding but requires training a 3D proposal network under supervision. (c): Our OV-SAM3D model effectively transfers the extensive knowledge embedded in the SAM to the 3D domain and enhances instance segmentation and recognition without requiring additional training. This capability enables open-vocabulary 3D scene understanding across various environments.
Given the difficulties and costs associated with collecting 3D data, transferring knowledge from foundation models in the 2D domain to the 3D domain presents a viable solution for achieving zero-shot learning in 3D scene understanding. Existing studies, such as SAM3D [42], SAMPro3D [41] and SAI3D [43], demonstrate effective 3D instance segmentation. However, these methods are constrained by the fact that SAM focuses solely on segmentation without incorporating recognition, thereby falling short of achieving full 3D scene understanding. On the basis of these zero-shot segmentation methods, advancing open-vocabulary 3D scene understanding requires the model to locate and recognize objects within a 3D scene through text guidance, regardless of whether the objects have been seen before. Related works, such as OpenScene [28], OpenMask3D [37] and Open3DIS [27], can achieve open-vocabulary 3D scene understanding when trained on datasets containing similar scenes. However, when the domain gap is substantial, nearly all of these methods fail, hindering the achievement of open-vocabulary 3D scene understanding in open-world scenarios. To address the challenge of understanding any 3D scene in an open-world, we propose a training-free framework that efficiently performs 3D scene understanding tasks without being constrained by domain gaps.
@@
In this work, we introduce OV-SAM3D, a universal open-vocabulary 3D scene understanding framework capable of performing understanding tasks on any given scene without prior knowledge about the scene. Inspired by the classic graph-based segmentation algorithm [13], our OV-SAM3D introduces superpoints as initial 3D prompts, which are subsequently merged into rough 3D masks. Then we calculate the designed overlapping scores through the back-projected segmentation results of SAM [20] for further refinement. By leveraging RAM’s recognition capabilities in posed images of the 3D scene and utilizing ChatGPT for screening, OV-SAM3D can generate open-world image tags matched with the initial 3D prompts using vision-language models like CLIP [29]. Based on the overlapping scores and matched tags, OV-SAM3D refines the rough 3D masks to produce final labeled 3D instances. In contrast to existing methods, our approach not only overcomes the limitations of class-agnostic masks in transferring SAM knowledge from 2D to 3D but also achieves superior results in open-vocabulary 3D scene understanding(Fig. 1). As a training-free 3D scene understanding algorithm, OV-SAM3D can comprehend arbitrary 3D scenes without prior knowledge, such as the categories of objects present and the consistent features of the domain.
@@
To compare with existing methods, we conduct experiments on the well-studied ScanNet200 dataset [33], but the closed-set environment goes against our original intention of understanding any 3D scene in an unknown open-world. Consequently, we also performed experiments on the outdoor nuScenes dataset [4], where few existing algorithms have been applied. The results show that OV-SAM3D is superior to existing state-of-the-art methods in an open-world context.
Figure 2: Overview of our OV-SAM3D that consists of two sub-modules: 1) SAM-centric Coarse Mask Generation first adapts a graph-based over-segmentation method to generate superpoints and selects some as initial 3D prompts to guide SAM. Then we revise the initial 3D prompts through the masks of SAM and create an overlapping score table. 2) Open Tags Guided Coarse Mask Merging combines the overlapping score table and reasonable open tags recognized by RAM, thus we can achieve the open-vocabulary 3D scene understanding task to get 3D instances with labels.
@@
  • We present OV-SAM3D, a universal framework for open-vocabulary 3D scene understanding, capable of understanding any 3D scenes without prior knowledge.
@@ 2 Related Work
Multimodal Foundation Models. The foundational models have gradually solidified their critical importance in the broader domain of artificial intelligence [35, 34, 12]. In particular, multimodal foundational models, which integrate data from various modalities, have emerged as a prominent area of research [29, 14, 11, 45]. CLIP [29]conducts contrastive learning on large datasets, enabling the model to classify various images effectively. Based on this, MaskCLIP [47] enhances the image-text classification capabilities by employing feature upsampling, extending functionality to coarse image segmentation. Tag2Text [18] introduces a foundational image tagging model that generates a broader range of tag categories and embeds these image tags into the visual-language model to guide downstream tasks. RAM [46] further innovates by combining tagging and captioning tasks, thereby expanding the recognition capabilities of the foundational image tagging model. As a landmark foundation model in the vision community, SAM [20] can generate high-quality masks through point prompts, box prompts, and text prompts. Through deep integration of SAM and CLIP, OV-SAM [45] can generate high-quality semantic segmentation masks for any image. Our research attempts to effectively combine the recognition capability of RAM, the localization strengths of SAM, and the classification power of CLIP to accomplish comprehensive 3D scene understanding tasks.
@@
Zero-Shot 3D Scene Segmentation. The impressive performance of SAM in zero-shot image segmentation tasks has greatly advanced the development of foundation models in computer vision. Its derivative models can be applied to challenging fields of segmentation such as medical images and remote sensing images and even extended to 3D point clouds for zero-shot 3D segmentation [6, 7, 32, 25, 5, 16]. Recent research increasingly relies on SAM to implement foundational model functions in the 3D domain due to the high cost of collecting extensive 3D point cloud data. SAM3D [42] projects multiple segmentation masks generated by SAM’s automatic segmentation onto a 3D scene, where adjacent point clouds are iteratively merged to produce a comprehensive 3D mask that encompasses the entire scene. Following the preliminary attempts with SAM3D, SAMPro3D [41] refines the segmentation process by sampling points from the 3D scene and projecting them into the posed images as point prompts to aid SAM in achieving more accurate segmentation results. SAI3D [43] introduces over-segmented results, which are then merged using affinity to generate the final 3D segmentation masks. However, these zero-shot 3D scene segmentation methods inherit the class-agnostic nature from SAM, making them unable to recognize specific segmented instances. In contrast to previous methods, our research does not rely on generating class-agnostic 3D masks followed by combining them with models like CLIP [29] or GroundingDINO [23] to obtain labels. Instead, we embed relevant semantic information directly into the mask generation process, enabling us to better address the challenges of understanding any 3D scene.
@@
Open-Vocabulary 3D Scene Understanding. Given any ambient text queries, open-vocabulary 3D scene understanding tasks [10, 17] aims to locate and identify matching instances, regardless of whether the model has previously encountered these instances. OpenScene [28] leverages the frozen 2D scene understanding model [15, 22] to train an aligned 3D encoder for extraction of dense features from point clouds in a 3D scene, and then ensembles the 2D&3D features o closely match arbitrary text queries. OpenIns3D [19] first employs a trained 3D network to generate class-agnostic mask proposals, which are then projected onto synthetic scene-level images [2, 21]. With the 2D scene understanding model detecting the position and category of objects in these scene-level images, OpenIns3D matches the projected mask proposals to achieve open-vocabulary 3D scene understanding. OpenMask3D [37] also utilizes Mask3D [36] to generate class-agnostic masks for 3D scenes, which are projected onto the posed images as point prompts. The method then extracts average CLIP embeddings from multi-scale image crops derived from SAM’s segmentation results, matching these embeddings with the query texts to segment 3D instances. Compared to previous studies, the significant improvement of our method is that we abandon training-required 3D encoders, which limit the generalization ability of models, and explore the collaborative effects of foundation models to achieve a universal open-vocabulary 3D scene understanding algorithm.
@@ 3 Methodology
Fig. 2 shows the detailed framework of our OV-SAM3D model. Given any scene and its associated set of posed RGB-D images, we draw inspiration from recent works [42, 43] to generate approximately over-segmented superpoints. The core of OV-SAM3D lies in merging these superpoints into accurate 3D instances and obtaining corresponding labels throughout the process. Firstly, we select high-quality superpoints as the initial 3D prompts, which are then projected onto posed images as point prompts to guide SAM in fine segmentation. The segmented mask is subsequently projected back into the 3D scene to refine the initial 3D prompts into coarse masks, from which overlapping scores are extracted. Concurrently, we utilize RAM [46] to identify open tags from the corresponding frames and employ ChatGPT to filter out unreasonable tags. By combining overlapping scores with matched open tags using CLIP, we obtain the final 3D instances and their labels.
@@ 3.1 SAM-centric Coarse Mask Generation
Directly processing unstructured 3D point clouds in 3D scenes significantly impacts computational efficiency, while grouping 3D points with similar geometric properties into continuous regions as superpoints remarkably accelerates the processing of OV-SAM3D. Drawing on the graph based image segmentation algorithm [13], we generate superpoints $\{S_{m}\in\mathbb{R}^{P_{m}\times 3}\}_{m=1}^{M}$ and select high-quality superpoints as the initial 3D prompts $\{Q_{n}\}_{n=1}^{N}\subseteq S$ based on $P_{m}$, which denotes the number of points in superpoint $S_{m}$. We then project the 3D prompts onto the top $T$ views with the lowest occlusion levels, assessed by the number of visible points in the depth pictures, and sample the projected pixel points as point prompts to guide SAM in fine segmentation.
@@
Considering the initial 3D prompts as the origins of the final 3D instance, we design an overlapping score table $F\in\mathbb{R}^{M\times N}$ to determine the association of all superpoints with their respective origins. We then back-project the $n$-th 2D mask onto the $t$-th view as $BP_{nt}\in\mathbb{R}^{P_{nt}\times 3},n\leq N,t\leq T$, to revise the randomly segmented 3D mask into a coarse 3D mask, where $P_{nt}$ denotes the number of points included in the 3D back-projection.
@@
The basis of this revision process and subsequent merging procedure is the designed overlapping score table $F$, generated as outlined in Algorithm 1. Firstly, we calculate the overlap region of all superpoints with the back-projection in the same 3D scene. Then, we update the score at the corresponding position $F_{ij}$, based on whether the overlap region of $\{S_{m}\cap Q_{n}\}$ meets a certain threshold $\theta$. After merging all superpoints into their respective 3D prompts, we obtain the overlapping score table $F$, as visualized in Fig. 3. In this table, along the row axis, the maximum index in each row $F_{m}^{1\times N}$ indicates the 3D prompt to which superpoint $S_{m}$ should belong, allowing us to merge all superpoints into the corresponding 3D prompts, resulting in coarse 3D masks $R_{c}\in\mathbb{R}^{P\times N}$, where $P$ denotes the total number of points in the scene. Along the column axis, each column $F_{n}^{M\times 1}$ represents the overlap distribution between each coarse 3D mask and all superpoints, which we define as the overlapping score of that coarse 3D mask.
Figure 3: Visualization of overlapping scores. We calculate the overlapping scores between queried coarse mask with other all other coarse masks. If the score is higher than a certain threshold, it indicates that these two coarse masks belong to the same instance. Here, we visualize some overlapping scores from high to low.
@@ 3.2 Open Tags Guided Coarse Mask Merging
To generate final high-quality 3D instance masks, it is necessary to remerge the initial coarse 3D masks due to the uncertainty associated with the initial 3D prompts in any given scene. This remerging process is determined by whether the total score $TS_{n}\in\mathbb{R}^{M\times 1}$, representing the overlap between coarse 3D masks, exceeds a certain threshold. This threshold is derived from the overlap score itself and dynamic parameter selected by the user.
@@
where $TS_{n}$ is the similarity score of $n$-th coarse 3D mask with other masks. $\tau$ is a parameter that can be modified by user-interaction. Note that OV-SAM3D provides several user-adjustable hyperparameters, with only their ranges constrained. This allows users to better adapt the model for any 3D scene in an open-world environment.
Limited by the narrow perspective of posed images, some large instances may not have overlapping regions across different views, making it impossible to merge them directly based on overlapping scores. As shown in Fig. 4, we employ an updatable merging strategy that continuously updates the coarse 3D mask to reflect the merged result while recording the composition of the merged coarse 3D masks throughout the process. This approach ultimately enables the generation of 3D instances along with detailed composition records.
@@
OV-SAM3D implements a universal open 3D scene understanding framework with the strict requirement that the model cannot acquire any prior knowledge of the 3D scenes it is intended to understand. Without prior knowledge of potential objects in a 3D scene, we introduce RAM to perform preliminary recognition on the corresponding posed images of the scene, as shown in Fig. 5. RAM provides a rough identification of which instances or labels are present. However, the open tag library of RAM includes not only instance tags but also tags representing colors, material properties, room types, and other instance-independent tags that are not conducive to 3D scene understanding tasks. To address this, we employ ChatGPT as a filter following RAM to select only the instance tags recognized by RAM, thereby generating the final open instance tags.
@@
After merging coarse masks into 3D instances and recognizing open instance tags, we match the tags with the 3D instance segmentation masks through CLIP to achieve the open-vocabulary 3D scene understanding task. To maintain matching accuracy without increasing algorithmic complexity, we avoid reprojecting the 3D masks onto posed images. Instead, we extract image crop features from existing SAM-segmented 2D masks using CLIP’s image encoder and match the CLIP text features $FT\in\mathbb{R}^{C\times 768}$ extracted from open tags with the image crop features $\{FC_{k}\in\mathbb{R}^{I_{k}\times 768}\}_{k=1}^{K}$, where $I_{k}$ represents the number of coarse masks in the $k$-th instance according to the composition records mentioned before:
@@
where $CS_{ki}$ represents the cosine similarity between the $i$-th course mask in the $k$-th instance and each of the $C$ text queries. For each coarse mask, we can choose the tag with the highest similarity as the label $\{Label_{i}\}_{i=1}^{I_{k}}$, but the matching result cannot be directly used as the label of the final 3D instance. We list the coarse 3D masks within each 3D instance based on the composition records and align the open tags $Label_{i}$ with the highest feature similarity $CS_{ki}$ as the final label $Label_{k}$ for the corresponding 3D segmentation instance. Subsequent ablation study in Tab. 3 demonstrates the effectiveness of the strategy of selecting the open-label with the highest feature similarity as the final 3D instance label.